# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ 3.2 Latent Diffusion Modeling
$\mathcal{L}(\mathbf{s})$ $=\frac{1}{T}\int_{0}^{T}\left(\mathbb{E}_{{\widehat{p}}^{*}_{data}(\mathbf{x})}\mathbb{E}_{\mathcal{N}(\mathbf{x}_{t};\mathbf{x},(1-t)\sigma^{2}\mathbf{I})}\left\|{\mathbf{s}(t,\mathbf{x}_{t})+\frac{\mathbf{x}_{t}-\mathbf{x}}{(1-t)\sigma^{2}}}\right\|^{2}\right)\mathrm{d}t$
$=\frac{1}{T}\int_{0}^{T}\left(\mathbb{E}_{{\widehat{p}}^{*}_{data}(\mathbf{x})}\mathbb{E}_{\mathcal{N}(\mathbf{z};\mathbf{0},\mathbf{I})}\left\|{\mathbf{s}(t,\mathbf{x}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sqrt{1-t}\sigma}}\right\|^{2}\right)\mathrm{d}t.$
Given $n$ i.i.d. samples $\{\widehat{\mathbf{x}}_{i}\}_{i=1}^{n}$ from ${\widehat{p}}^{*}_{data}(\cdot)$, and $m$ i.i.d. samples $\{(t_{j},\mathbf{z}_{j})\}_{j=1}^{m}$ from $\text{U}[0,T]$ and $\mathcal{N}(\mathbf{z};\mathbf{0},\mathbf{I}_{d^{*}})$, we can employ the empirical risk minimizer (ERM) to estimate the score $\mathbf{s}^{*}$. This ERM, denoted as $\widehat{\mathbf{s}}$, is determined by
$\widehat{\mathbf{s}}\in{\mathop{\rm{argmin}}_{\mathbf{s}\in\text{NN}}}\leavevmode\nobreak\ \widehat{\mathcal{L}}(\mathbf{s}):=\frac{1}{mn}\sum_{j=1}^{m}\sum_{i=1}^{n}\left\|{\mathbf{s}(t_{j},\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t_{j}}\mathbf{z}_{j})+\frac{\mathbf{z}_{j}}{\sigma\sqrt{1-t_{j}}}}\right\|^{2},$ (6)
@@
We observe that the SDE represented by (7) is well-defined. More precisely, it possesses a unique strong solution attributed to the Lipschitz continuity of the score function $\widehat{\mathbf{s}}(t,\mathbf{x})$ with respect to $\mathbf{x}$, as substantiated in (11). Now, we can employ a discrete-time approximation for the sampling dynamics (7). Let
$0=t_{0}<t_{1}<\cdots<t_{N}=T,\leavevmode\nobreak\ N\in\mathbb{N}^{+},$
@@
$\mathrm{d}\widetilde{\mathbf{x}}_{t}=\sigma^{2}\widehat{\mathbf{s}}(t_{k},\widetilde{\mathbf{x}}_{t_{k}})\mathrm{d}t+\sigma\mathrm{d}\mathbf{w}_{t},\leavevmode\nobreak\ \widetilde{\mathbf{x}}_{0}=\mathbf{x}_{0}\sim{q}(\sigma,\mathbf{x}),\leavevmode\nobreak\ t\in[t_{k},t_{k+1}),$ (8)
for $k=0,1,\cdots,N-1$. Subsequently, we can utilize dynamics (8) to generate new samples by using the trained decoder in Section 3.1.
@@
  1. Input: $\sigma$, $N$, $d^{*}$, $\{\mathbf{x}_{i}\}_{i=1}^{n}\sim p_{data}$, $\{\mathbf{y}_{i}\}_{i=1}^{\mathcal{M}}\sim\tilde{p}_{data}$, $\{(t_{j},\mathbf{z}_{j})\}_{j=1}^{m}\sim\text{U}[0,T]\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \mathcal{N}(\mathbf{z};\mathbf{0},\mathbf{I}_{d^{*}})$.
@@ 4. Sampling procedure:
  for all $k=0,1,\ldots,N-1$ do
     Sample $\mathbf{\epsilon}_{k}\sim{\mathcal{N}}(\mathbf{0},\mathbf{I}_{d^{*}})$;
     $\mathbf{b}(\widetilde{\mathbf{x}}_{k})=\widehat{\mathbf{s}}\left(\sqrt{1-\frac{k}{N}}\sigma,\widetilde{\mathbf{x}}_{k}\right)$;
     $\widetilde{\mathbf{x}}_{k+1}=\widetilde{\mathbf{x}}_{k}+\frac{\sigma^{2}}{N}\mathbf{b}(\widetilde{\mathbf{x}}_{k})+\frac{\sigma}{N}\mathbf{\epsilon}_{k}$;
@@
  5. Output: $\widehat{\boldsymbol{D}}(\widetilde{\mathbf{x}}_{N})$.
@@ 4.1 Error Bounds in Pre-training
The pre-trained data distribution $\widetilde{p}_{data}$ is supported on $[0,1]^{d}$.
@@
There exist continuously differential functions $\boldsymbol{E}^{*}:[0,1]^{d}\rightarrow[0,1]^{d^{*}}$ and $\boldsymbol{D}^{*}:[0,1]^{d^{*}}\rightarrow\mathbb{R}^{d}$ such that $\mathcal{H}(\boldsymbol{E},\boldsymbol{D})$ attains its minimum. The minimum value is denoted by $\varepsilon_{\mathrm{compress}-{\mathrm{noise}}}:=\mathcal{H}(\boldsymbol{E}^{*},\boldsymbol{D}^{*})\leq\delta_{0}$, where $\delta_{0}\geq 0$. Furthermore, $\boldsymbol{E}^{*}\in\mathrm{Lip}(\xi_{\boldsymbol{E}}),\boldsymbol{D}^{*}\in\mathrm{Lip}(\xi_{\boldsymbol{D}})$ respectively.
@@
When the data is perfectly compressible, implying the existence of $\boldsymbol{D}^{*},\boldsymbol{E}^{*}$ such that $\boldsymbol{D}^{*}\circ\boldsymbol{E}^{*}=\mathbf{I}_{d}$, then $\delta_{0}=0.$
@@
Suppose that Assumptions 4.1-4.2 hold, then we have
$\mathbb{E}_{\mathcal{Y}}\mathbb{E}_{\mathbf{y}\sim\widetilde{p}_{data}}\|(\widehat{\boldsymbol{D}}\circ\widehat{\boldsymbol{E}})(\mathbf{y})-\mathbf{y}\|^{2}=\mathcal{O}({\mathcal{M}}^{-1/(d+2)}+\delta_{0}).$
@@ 4.2 Error Bounds in Latent Diffusion Modeling
To obtain error bounds in latent diffusion modeling, we systematically deconstruct the error term into three distinct components, as expressed in the following inequality:
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widetilde{\pi}_{T}},\widehat{p}^{*}_{data})]\leq\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widehat{\pi}_{T}},\pi_{T})]+\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\pi}_{T},{\widetilde{\pi}_{T}})]+\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},\widehat{p}^{*}_{data})].$ (9)
Here, $\pi_{t}$, $\widehat{\pi}_{t},\widetilde{\pi}_{t}$ denote the distributions of $\mathbf{x}_{t},\widehat{\mathbf{x}}_{t},\widetilde{\mathbf{x}}_{t}$ correspondingly. Subsequently, we proceed to establish individual bounds for the three terms on the right-hand side of (9) in Sections 4.2.1-4.2.3. The integration of these three analyses culminates in the oracle inequality for $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widetilde{\pi}_{T}},p_{data})]$, which is expounded upon in Section 4.2.4. To facilitate this analysis, we first introduce two necessary assumptions.
Assumption 4.3 (Bounded Support).
The target distribution $p_{data}$ is supported on $[0,1]^{d}$.
Assumption 4.4.
The distribution $\widehat{p}^{*}_{data}$ in the latent space is twice continuously differentiable and satisfies $-\alpha\mathrm{I}_{d^{*}}\preccurlyeq\nabla^{2}\log\widehat{p}^{*}_{data}\preccurlyeq\alpha\mathrm{I}_{d^{*}}$ with $\alpha\sigma^{2}>1$.
These two assumptions assume a paramount role in providing a solid underpinning for the following analytical frameworks. The bounded support assumption (Assumption 4.3), is also elucidated in [LLT23, Assumption 2], [LWCC23], and [OAS23, Assumption 2.4], emerging as a foundational pillar in the scrutiny of diffusion models. Shifting our focus to the smoothness assumption (Assumption 4.4), a meticulous examination unveils a nuanced exploration across multiple works. Contributions from [LLT22, Assumption 1], [CLL23, Assumptions 3-4], [LLT23, Assumptions 3-4], [GNZ23, Assumption 1], [OAS23, Assumptions 2.4, 2.6], and [CHZW23, Assumption 3] collectively contribute to a multifaceted comprehension of this smoothness assumption. Moreover, [CHZW23, Assumptions 1, 2] introduced analogous conditions. Their Assumption 1 delineates the representation of data points $\mathbf{x}\sim p_{data}$ as a product of an unknown matrix $A$ with orthonormal columns and a latent variable $\mathbf{z}$ following a distribution $P_{z}$ with a density function $p_{z}$. Meanwhile, their Assumption 2 supposes the density function $p_{z}>0$ is twice continuously differentiable. Moreover, there exist positive constants $B,C_{1},C_{2}$ such that when $\|\mathbf{z}\|_{2}\geq B$, the density function $p_{z}(\mathbf{z})\leq$ $(2\pi)^{-d/2}C_{1}\exp\left(-C_{2}\|\mathbf{z}\|_{2}^{2}/2\right)$.
Under Assumptions 4.3-4.4, we can further establish the Lipschitz continuity property for the convoluted score function, as demonstrated in the following two lemmas.
Lemma 4.2.
Suppose that Assumption 4.3 holds and let $T\in(0,1)$. Then $\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$ is $\xi$-Lipschitz continuous w.r.t. $\mathbf{x}$ on $[0,T]\times\mathbb{R}^{d^{*}}$, where $\xi\leq{\rm{max}}\left\{\frac{d^{*}}{\sigma^{4}(1-T)^{2}},\frac{1}{\sigma^{2}(1-T)}\right\}$.
Lemma 4.3.
Suppose that Assumption 4.4 holds and let $T\in(0,1)$. Then $\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$ is $\xi$-Lipschitz continuous w.r.t. $\mathbf{x}$ on $[0,T]\times\mathbb{R}^{d^{*}}$, where $\xi=\alpha(1+\alpha d^{*})$.
4.2.1 Bound $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widehat{\pi}_{T}},\pi_{T})]$
In this section, our aim is to establish an upper bound for $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widehat{\pi}_{T}},\pi_{T})]$. Employing coupling techniques, it follows that this error emanates from the score estimation. Consequently, we commence by obtaining an error bound for the estimated score $\widehat{\mathbf{s}}$ in Theorem 4.6. Following this, we proceed to derive the upper bound for $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widehat{\pi}_{T}},\pi_{T})]$ as delineated in Theorem 4.7.
@@ 4.2.1 Bound $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widehat{\pi}_{T}},\pi_{T})]$
$\ell_{\mathbf{s}}(\widehat{\mathbf{x}}):=\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}\left\|\mathbf{s}(t,\mathbf{x}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sigma\sqrt{1-t}}\right\|^{2}\mathrm{d}t,$
@@
^$\widehat{\ell}_{\mathbf{s}}(\widehat{\mathbf{x}}):=\frac{1}{m}\sum_{j=1}^{m}\left\|\mathbf{s}(t_{j},\widehat{\mathbf{x}}+\sigma\sqrt{1-t_{j}}\mathbf{z}_{j})+\frac{\mathbf{z}_{j}}{\sigma\sqrt{1-t_{j}}}\right\|^{2}.$
@@
$\mathcal{L}(\widehat{\mathbf{s}})-\mathcal{L}(\mathbf{s}^{*})$ $=\mathcal{L}(\widehat{\mathbf{s}})-2\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})+\mathcal{L}(\mathbf{s}^{*})+2\left(\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})-\mathcal{L}(\mathbf{s}^{*})\right)$
@@
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}}\|\widehat{\mathbf{s}}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}\mathrm{d}t\right)$
@@
$\leq$ $\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathcal{L}(\widehat{\mathbf{s}})-2\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})+\mathcal{L}(\mathbf{s}^{*})\right)+2\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})-\widehat{\mathcal{L}}_{\widehat{\mathcal{X}},\mathbb{T},\mathbb{Z}}(\widehat{\mathbf{s}})\right)+2\mathop{\mathrm{inf}}_{\mathbf{s}\in{\mathrm{NN}}}(\mathcal{L}(\mathbf{s})-\mathcal{L}(\mathbf{s}^{*})).$
@@
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathcal{L}(\widehat{\mathbf{s}})-2\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})+\mathcal{L}(\mathbf{s}^{*})\right)+2\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})-\widehat{\mathcal{L}}_{\widehat{\mathcal{X}},\mathbb{T},\mathbb{Z}}(\widehat{\mathbf{s}})\right)$

### v2

@@ 3.2 Latent Diffusion Modeling
$\mathcal{L}(\mathbf{s})$ $=\frac{1}{T}\int_{0}^{T}\left(\mathbb{E}_{{\widehat{p}}^{*}_{data}(\mathbf{x})}\mathbb{E}_{\mathcal{N}(\mathbf{x}_{t};\mathbf{x},(1-t)\sigma^{2}\mathbf{I})}\left\|{\mathbf{s}(1-t,\mathbf{x}_{t})+\frac{\mathbf{x}_{t}-\mathbf{x}}{(1-t)\sigma^{2}}}\right\|^{2}\right)\mathrm{d}t$
$=\frac{1}{T}\int_{0}^{T}\left(\mathbb{E}_{{\widehat{p}}^{*}_{data}(\mathbf{x})}\mathbb{E}_{\mathcal{N}(\mathbf{z};\mathbf{0},\mathbf{I})}\left\|{\mathbf{s}(1-t,\mathbf{x}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sqrt{1-t}\sigma}}\right\|^{2}\right)\mathrm{d}t.$
Given $n$ i.i.d. samples $\{\widehat{\mathbf{x}}_{i}\}_{i=1}^{n}$ from ${\widehat{p}}^{*}_{data}(\cdot)$, and $m$ i.i.d. samples $\{(t_{j},\mathbf{z}_{j})\}_{j=1}^{m}$ from $\mathrm{U}[0,T]$ and $\mathcal{N}(\mathbf{z};\mathbf{0},\mathbf{I}_{d^{*}})$, we can employ the empirical risk minimizer (ERM) to estimate the score $\mathbf{s}^{*}$. This ERM, denoted as $\widehat{\mathbf{s}}$, is determined by
$\widehat{\mathbf{s}}\in{\mathop{\rm{argmin}}_{\mathbf{s}\in\mathrm{NN}}}\leavevmode\nobreak\ \widehat{\mathcal{L}}(\mathbf{s}):=\frac{1}{mn}\sum_{j=1}^{m}\sum_{i=1}^{n}\left\|{\mathbf{s}(1-t_{j},\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t_{j}}\mathbf{z}_{j})+\frac{\mathbf{z}_{j}}{\sigma\sqrt{1-t_{j}}}}\right\|^{2},$ (6)
@@
Now, we employ a discrete-time approximation for the sampling dynamics (7). Let
$0=t_{0}<t_{1}<\cdots<t_{K}=T,\leavevmode\nobreak\ K\in\mathbb{N}^{+},$
@@
$\mathrm{d}\widetilde{\mathbf{x}}_{t}=\sigma^{2}\widehat{\mathbf{s}}(t_{i},\widetilde{\mathbf{x}}_{t_{i}})\mathrm{d}t+\sigma\mathrm{d}\mathbf{w}_{t},\leavevmode\nobreak\ \widetilde{\mathbf{x}}_{0}=\mathbf{x}_{0}\sim{q}(\sigma,\mathbf{x}),\leavevmode\nobreak\ t\in[t_{i},t_{i+1}),$ (8)
for $i=0,1,\cdots,K-1$. Subsequently, we can utilize dynamics (8) to generate new samples by using the trained decoder in Section 3.1.
@@
  1. Input: $\sigma$, $N$, $d^{*}$, $\{\mathbf{x}_{i}\}_{i=1}^{n}\sim p_{data}$, $\{\mathbf{y}_{i}\}_{i=1}^{\mathcal{M}}\sim\widetilde{p}_{data}$, $\{(t_{j},\mathbf{z}_{j})\}_{j=1}^{m}\sim\text{U}[0,T]\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \mathcal{N}(\mathbf{z};\mathbf{0},\mathbf{I}_{d^{*}})$.
@@ 4. Sampling procedure:
  for all $i=0,1,\ldots,K-1$ do
     Sample $\mathbf{\epsilon}_{i}\sim{\mathcal{N}}(\mathbf{0},\mathbf{I}_{d^{*}})$;
     $\mathbf{b}(\widetilde{\mathbf{x}}_{i})=\widehat{\mathbf{s}}\left(\sqrt{1-\frac{i}{K}}\sigma,\widetilde{\mathbf{x}}_{i}\right)$;
     $\widetilde{\mathbf{x}}_{i+1}=\widetilde{\mathbf{x}}_{i}+\frac{\sigma^{2}}{K}\mathbf{b}(\widetilde{\mathbf{x}}_{i})+\frac{\sigma}{K}\mathbf{\epsilon}_{i}$;
@@
  5. Output: $\widehat{\boldsymbol{D}}(\widetilde{\mathbf{x}}_{K})$.
@@ 4.1 Error Bounds in Pre-training
The pre-trained data distribution $\widetilde{p}_{data}$ is supported on $[-1,1]^{d}$.
@@
There exist continuously differential functions $\boldsymbol{E}^{*}:[-1,1]^{d}\rightarrow[-1,1]^{d^{*}}$ and $\boldsymbol{D}^{*}:[-1,1]^{d^{*}}\rightarrow\mathbb{R}^{d}$ such that $\mathcal{H}(\boldsymbol{E},\boldsymbol{D})$ attains its minimum. The minimum value is denoted by $\varepsilon_{\mathrm{compress}-{\mathrm{noise}}}:=\mathcal{H}(\boldsymbol{E}^{*},\boldsymbol{D}^{*})\leq\delta_{0}$, where $\delta_{0}\geq 0$. Furthermore, $\boldsymbol{E}^{*}\in\mathrm{Lip}(\xi_{\boldsymbol{E}}),\boldsymbol{D}^{*}\in\mathrm{Lip}(\xi_{\boldsymbol{D}})$ respectively.
@@
These two assumptions assume a paramount role in providing a solid underpinning for the pre-training framework. Additionally, the theoretical analysis of diffusion models elucidates the bounded support assumption on the target distribution, as presented in [LLT23, Assumption 2], [LWCC23], and [OAS23, Assumption 2.4], establishing it as a foundational pillar in the study of diffusion models. In Assumption 4.2, when the data is perfectly compressible, implying the existence of $\boldsymbol{D}^{*},\boldsymbol{E}^{*}$ such that $\boldsymbol{D}^{*}\circ\boldsymbol{E}^{*}=\mathbf{I}_{d}$, then $\delta_{0}=0.$
Lemma 4.1.
Suppose that Assumptions 4.1-4.2 hold. Then, for any $\epsilon>0$, there exist two ReLU neural networks $\boldsymbol{E}\in\mathcal{E}$ and $\boldsymbol{D}\in\mathcal{D}$ with respective parameters
$L_{\boldsymbol{E}}=\mathcal{O}\left(\log\frac{1}{\epsilon}+d\right),M_{\boldsymbol{E}}=\mathcal{O}\left(d^{*}\xi_{\boldsymbol{E}}^{d}\epsilon^{-d}\right),J_{\boldsymbol{E}}=\mathcal{O}\left(d^{*}\xi_{\boldsymbol{E}}^{d}\epsilon^{-d}\left(\log\frac{1}{\epsilon}+d\right)\right),\kappa_{\boldsymbol{E}}=\mathcal{O}\left(\max\{1,\xi_{\boldsymbol{E}}\}\right),$
$L_{\boldsymbol{D}}=\mathcal{O}\left(\log\frac{1}{\epsilon}+d^{*}\right),M_{\boldsymbol{D}}=\mathcal{O}\left(d\xi_{\boldsymbol{D}}^{d^{*}}\epsilon^{-d^{*}}\right),J_{\boldsymbol{D}}=\mathcal{O}\left(d\xi_{\boldsymbol{D}}^{d^{*}}\epsilon^{-d^{*}}\left(\log\frac{1}{\epsilon}+d^{*}\right)\right),\kappa_{\boldsymbol{D}}=\mathcal{O}\left(\max\{1,\xi_{\boldsymbol{D}}\}\right),$
such that
$\mathcal{H}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E}^{*},\boldsymbol{D}^{*})\lesssim\epsilon.$
Moreover, $\boldsymbol{E}$ and $\boldsymbol{D}$ are Lipschitz continuous with Lipschitz constants $\gamma_{\boldsymbol{E}}:=10d\xi_{\boldsymbol{E}}$ and $\gamma_{\boldsymbol{D}}:=10d^{*}\xi_{\boldsymbol{D}}$, respectively. That is, for any $\mathbf{y}_{1},\mathbf{y}_{2}\in[-1,1]^{d}$, $\boldsymbol{E}$ satisfies
$\|\boldsymbol{E}(\mathbf{y}_{1})-\boldsymbol{E}(\mathbf{y}_{2})\|_{\infty}\leq\gamma_{\boldsymbol{E}}\|\mathbf{y}_{1}-\mathbf{y}_{2}\|,$
for any $\mathbf{y}_{1},\mathbf{y}_{2}\in[-1,1]^{d^{*}}$, $\boldsymbol{D}$ satisfies
And there exist two constants $K_{\boldsymbol{D}}=\mathcal{O}(1)$, $K_{\boldsymbol{E}}=\mathcal{O}(1)$ such that
$\sup_{\mathbf{y}\in[-1,1]^{d}}\boldsymbol{E}(\mathbf{y})\leq K_{\boldsymbol{E}},\leavevmode\nobreak\ \sup_{\mathbf{y}\in[-1,1]^{d^{*}}}\boldsymbol{D}(\mathbf{y})\leq K_{\boldsymbol{D}}.$
Now, we restrict $\boldsymbol{E}\in\mathcal{E}$ and $\boldsymbol{D}\in\mathcal{D}$ to the following two classes:
$\mathcal{E}_{0}:=\Big{\{}\boldsymbol{E}\in\mathcal{E}\Big{|}$ The network configurations satisfy
$L_{\boldsymbol{E}}=\mathcal{O}\left(\log\frac{1}{\epsilon}+d\right),M_{\boldsymbol{E}}=\mathcal{O}\left(d^{*}\xi_{\boldsymbol{E}}^{d}\epsilon^{-d}\right),$
$J_{\boldsymbol{E}}=\mathcal{O}\left(d^{*}\xi_{\boldsymbol{E}}^{d}\epsilon^{-d}\left(\log\frac{1}{\epsilon}+d\right)\right),\kappa_{\boldsymbol{E}}=\mathcal{O}\left(\max\{1,\xi_{\boldsymbol{E}}\}\right),\sup_{\mathbf{y}\in[-1,1]^{d}}\|\boldsymbol{E}(\mathbf{y})\|\leq K_{\boldsymbol{E}},$
$\forall\mathbf{y}_{1},\mathbf{y}_{2}\in[-1,1]^{d},\|\boldsymbol{E}(\mathbf{y}_{1})-\boldsymbol{E}(\mathbf{y}_{2})\|_{\infty}\leq\gamma_{\boldsymbol{E}}\|\mathbf{y}_{1}-\mathbf{y}_{2}\|\Big{\}},$
$\mathcal{D}_{0}:=\Big{\{}\boldsymbol{D}\in\mathcal{D}\Big{|}$ The network configurations satisfy
$L_{\boldsymbol{D}}=\mathcal{O}\left(\log\frac{1}{\epsilon}+d^{*}\right),M_{\boldsymbol{D}}=\mathcal{O}\left(d\xi_{\boldsymbol{D}}^{d^{*}}\epsilon^{-d^{*}}\right),$
$J_{\boldsymbol{D}}=\mathcal{O}\left(d\xi_{\boldsymbol{D}}^{d^{*}}\epsilon^{-d^{*}}\left(\log\frac{1}{\epsilon}+d^{*}\right)\right),\kappa_{\boldsymbol{D}}=\mathcal{O}\left(\max\{1,\xi_{\boldsymbol{D}}\}\right),\sup_{\mathbf{y}\in[-1,1]^{d^{*}}}\|\boldsymbol{D}(\mathbf{y})\|\leq K_{\boldsymbol{D}},$
$\forall\mathbf{y}_{1},\mathbf{y}_{2}\in[-1,1]^{d^{*}},\|\boldsymbol{D}(\mathbf{y}_{1})-\boldsymbol{D}(\mathbf{y}_{2})\|_{\infty}\leq\gamma_{\boldsymbol{D}}\|\mathbf{y}_{1}-\mathbf{y}_{2}\|\Big{\}}.$
Consequently, we can derive the error bound for the pre-training procedure, as shown in the following theorem.
@@
Suppose that Assumptions 4.1-4.2 hold. By taking $\epsilon=\mathcal{M}^{-\frac{1}{d+2}}$ in Lemma 4.1, the encoder estimator $\widehat{\boldsymbol{E}}$ with the neural network structures belonging to class $\mathcal{E}_{0}$ and the decoder estimator $\widehat{\boldsymbol{D}}$ with neural network structures belonging to class $\mathcal{D}_{0}$ satisfy
$\mathbb{E}_{\mathcal{Y}}\mathbb{E}_{\mathbf{y}\sim\widetilde{p}_{data}}\|(\widehat{\boldsymbol{D}}\circ\widehat{\boldsymbol{E}})(\mathbf{y})-\mathbf{y}\|^{2}=\widetilde{\mathcal{O}}({\mathcal{M}}^{-\frac{1}{d+2}}+\delta_{0}).$
@@ 4.2 Error Bounds in Latent Diffusion Modeling
In this section, we obtain the error bounds in latent diffusion modeling. We denote $\pi_{t}$, $\widehat{\pi}_{t}$, and $\widetilde{\pi}_{t}$ as the distributions of $\mathbf{x}_{t}$, $\widehat{\mathbf{x}}_{t}$, and $\widetilde{\mathbf{x}}_{t}$ correspondingly. Since the diffusion process (5) converges to $\widehat{p}_{data}^{*}$ as $t\rightarrow 1$ and the domain is compact, we define a truncated estimator $\pi_{T}^{L}$ for $\widehat{p}_{data}^{*}$ as the distribution of $\mathbf{x}_{T}\cdot\mathbf{I}_{\{\|\mathbf{x}_{T}\|_{\infty}\leq L\}}$, for a large constant $L>0$ such that $\pi_{T}^{L}\approx\widehat{p}_{data}^{*}$. Similarly, we define $\widehat{\pi}_{T}^{L}$ and $\widetilde{\pi}_{T}^{L}$ as the distributions of $\widehat{\mathbf{x}}_{T}\cdot\mathbf{I}_{\{\|\widehat{\mathbf{x}}_{T}\|_{\infty}\leq L\}}$ and $\widetilde{\mathbf{x}}_{T}\cdot\mathbf{I}_{\{\|\widetilde{\mathbf{x}}_{T}\|_{\infty}\leq L\}}$, respectively. Here, we truncate the supports of the distributions $\pi_{T},\widehat{\pi}_{T}$, and $\widetilde{\pi}_{T}$ to ensure a bounded support for the distribution estimators, which is done for technical reasons.
@@
$\ell_{\mathbf{s}}(\widehat{\mathbf{x}}):=\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}\left\|\mathbf{s}(1-t,\mathbf{x}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sigma\sqrt{1-t}}\right\|^{2}\mathrm{d}t,$
@@
^$\widehat{\ell}_{\mathbf{s}}(\widehat{\mathbf{x}}):=\frac{1}{m}\sum_{j=1}^{m}\left\|\mathbf{s}(1-t_{j},\widehat{\mathbf{x}}+\sigma\sqrt{1-t_{j}}\mathbf{z}_{j})+\frac{\mathbf{z}_{j}}{\sigma\sqrt{1-t_{j}}}\right\|^{2}.$
@@
$\mathcal{L}(\widehat{\mathbf{s}})-\mathcal{L}(\mathbf{s}^{*})$
$=\mathcal{L}(\widehat{\mathbf{s}})-2\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})+\mathcal{L}(\mathbf{s}^{*})+2\left(\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})-\mathcal{L}(\mathbf{s}^{*})\right)$
@@
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}}\|\widehat{\mathbf{s}}(1-t,\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathrm{d}t\right)$
@@
$\leq\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathcal{L}(\widehat{\mathbf{s}})-2\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})+\mathcal{L}(\mathbf{s}^{*})\right)+2\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})-\widehat{\mathcal{L}}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}(\widehat{\mathbf{s}})\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +2\mathop{\mathrm{inf}}_{\mathbf{s}\in{\mathrm{NN}}}(\mathcal{L}(\mathbf{s})-\mathcal{L}(\mathbf{s}^{*})).$
@@
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathcal{L}(\widehat{\mathbf{s}})-2\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})+\mathcal{L}(\mathbf{s}^{*})\right)+2\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})-\widehat{\mathcal{L}}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}(\widehat{\mathbf{s}})\right)$